OpenStudy (anonymous):

Please explain how to find the vertex for this problem. f(x)=(x+2)^2+9

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms-thumb.png <--- can you see the vertex coordinates?

OpenStudy (jdoe0001):

\(\bf f(x)=(x+2)^2+9\implies f(x)=[x-({\color{blue}{ -2}})]^2{\color{purple}{ +9}}\)

OpenStudy (anonymous):

Yes, I can see them

OpenStudy (anonymous):

so is the vertex (2,-9)? Or (-2,9)?

OpenStudy (anonymous):

f(x)=(x+2)^2+9..Make this into a quadratic equation ax^2+bx+c..The vertex for the x corinate is determined by the formula x=-b/2a,

OpenStudy (anonymous):

Okay, thank you!